# A Little Depth Goes a Long Way: The Expressive Power of Log-Depth Transformers

### v1

@@ 4 Log Depth Enables Recognizing Regular Languages
One natural problem that constant-depth transformers cannot express is recognizing regular languages, which is closely related to state tracking (Liu et al., 2023; Merrill et al., 2024). Liu et al. (2023, Theorem 1) show how a log-depth transformer can recognize regular languages using a binary tree construction similar to associative scan (Hillis & Steele Jr, 1986). However, their result requires simplifying assumptions, removing residual connections from the transformer and assuming specific positional encodings. As discussed in Section 2.2, dealing with residual connections is particularly tricky in universal transformers, requiring proper memory management of cells in the residual stream so that outputs from the previous iteration of a layer interfere with a later iteration. Our result therefore refines that of Liu et al. (2023) to hold with a more general universal transformer model that uses residual connections and does not rely on specific positional encodings:
Theorem 4.1 (Regular Language Recognition).
@@
Proof.
Regular language recognition can be framed as multiplying a sequence of elements in the automaton’s transition monoid (Myhill, 1957; Thérien, 1981). It thus suffices to show how $n$ elements in a finite monoid can be multiplied with $\Theta(\log n)$ depth. We show how a log-depth universal transformer can implement the standard binary tree construction (Barrington & Thérien, 1988; Liu et al., 2023; Merrill et al., 2024) where each level multiplies two items, meaning the overall depth is $\Theta(\log\lvert w\rvert)$. We will represent a tree over the input tokens within the transformer. Each level of the tree will take 8 transformer layers. We define a notion of active tokens: at level 0, all tokens are active, and, at level $\ell$, tokens at $t\cdot 2^{\ell}-1$ for any $t$ will remain active, and all other tokens will be marked as inactive. As an invariant, active token $i=t\cdot 2^{\ell}-1$ in level $\ell$ will store a unit-norm vector $\delta^{\ell}_{i}$ that represents the cumulative product of tokens from $i-2^{\ell}+1$ to $i$.
We now proceed by induction over $\ell$, defining the behavior of non-\$ tokens at layers that make up level $\ell$. The current group element $\delta^{\ell}_{i}$ stored at active token $i$ is, by inductive assumption, the cumulative product from $i-2^{\ell}+1$ to $i$. Let $\alpha^{\ell}_{i}$ denote that token $i$ is active. By Lem. A.6 we use a layer to store $i-1$ at token $i$. The next layer attends with query $\phi(i-1)$, key $\phi(j)$, and value $\delta^{\ell}_{j}$ to retrieve $\delta^{\ell}_{i-1}$, the group element stored at the previous token. Finally, another layer attends with query $\vec{1}$, key $\langle\phi(j)_{1},\alpha^{\ell}_{i}\rangle$, and value $\delta^{\ell}_{j-1}$ to retrieve the group element $\delta^{\ell}_{j^{*}}$ stored at the previous active token, which represents the cumulative product from $i-2\cdot 2^{\ell}+1$ to $i-2^{\ell}$. Next, we will use two layers to update $\delta^{\ell}_{i}\leftarrow\delta^{\ell+1}_{i}$ and $\delta^{\ell}_{j}\leftarrow\vec{0}$, which is achieved as follows. First, we assert there exists a single feedforward layer that uses a table lookup to compute $\delta^{\ell}_{j^{*}},\delta^{\ell}_{i}\mapsto d$ such that $d/\lVert d\rVert=\delta^{\ell}_{j^{*}}\cdot\delta^{\ell}_{i}=\delta^{\ell+1}_{i}$. Next, we invoke Lem. A.5 to construct a layer that adds $d$ to an empty cell of the residual stream and then another layer that deletes it. This second layer can now read both $\delta^{\ell}_{i},\delta^{\ell}_{j^{*}}$ and $\delta^{\ell+1}_{i}$ (from $d$) as input, and we modify it to add $\delta^{\ell+1}_{i}-\delta^{\ell}_{i}$ to $\delta^{\ell}_{i}$, changing its value to $\delta^{\ell+1}_{i}$. Similarly, we modify it to add $-\delta^{\ell}_{j^{*}}$ to $\delta^{\ell}_{j^{*}}$ to set it to 0. A feedforward network then subtracts $\delta^{\ell}_{i}$ from the residual stream and adds $\delta^{\ell}_{i}\cdot\delta^{\ell}_{j}$. This requires at most 4 layers.
To determine activeness in layer $\ell+1$, each token $i$ attends to its left to compute $c_{i}/i$, where $c_{i}$ is the prefix count of active tokens, inclusive of the current token. We then compute $\phi(c_{i}/i,1/i)=\phi(c_{i})$ and store $c_{i}$ it temporarily in the residual stream. At this point, we use Lem. 3.1 to construct 7 layers that compute $c_{i}\bmod 2$ with no storage overhead. The current token is marked as active in layer $\ell+1$ iff $c_{i}=0\mod 2$, which is equivalent to checking whether $i=t\cdot 2^{\ell}-1$ for some $t$. In addition to updating the new activeness $\alpha^{\ell+1}_{i}$, we also persist store the previous activeness $\alpha^{\ell}_{i}$ in a separate cell of the residual stream and clear $c_{i}$. This requires at most 8 layers.
Thm. 4.1 thus reveals that running a transformer to $\log n$ depth on inputs of length $n$ unlocks new power compared to a fixed-depth transformer. If we do not care that the construction is uniform across layers, we can simplify the block of 8 layers that determines activeness to 1 layer: we simply hardcode the layer index $\ell$ and use a single transformer layer to compute $i\mod\ell$. Thus, the non-uniform construction results in a family of shallower transformers:
@@
These results can be extended beyond regular languages: if a $b$-layer transformer can perform some binary associative operation $\oplus:X\times X\to X$, then one can construct an $\Theta(b\log n)$ layer transformer that computes the iterated version of the operator on $n$ values, $x_{1}\oplus x_{2}\oplus\ldots\oplus x_{n}\in X$. One natural iterated problem is iterated matrix multiplication. For matrices from a fixed set (e.g., $k\times k$ boolean matrices), Thm. 4.1 already shows that this task can be performed. However, if the matrices are not from a fixed set (e.g., matrices over $\mathbb{Z}$ or $\mathbb{Q}$ or whose shape depends on $n$), then it is unclear whether log-depth transformers can solve the binary multiplication problem, and thus whether they can solve the iterated version.
4.1 Fixed Depth and Bounded Length Inputs
Interestingly, while Thms. 4.1 and 4.2 are about log-depth transformers, they can be turned around to infer bounds on the input length up to which fixed depth transformers (i.e., depth fixed w.r.t. input length) can recognize regular languages. Specifically, given any regular language $L$ and a fixed $d$, Cor. 4.2 implies that there exists a depth $d$ transformer that can recognize strings $w\in L$ as long as $4\lceil\log_{2}\lvert w\rvert\rceil+5\leq d$ [^4] The inequality holds because Cor. 4.2 can be generalized to state that $T_{n}$ correctly recognizes all strings of length $m\leq n$.
@@ 4.1 Fixed Depth and Bounded Length Inputs
Let $L$ be a regular language over $\Sigma$ and \$$\$\not\in\Sigma$. For any $d\in\mathbb{N}$, there exists a causally masked $d$-layer transformer that, on any string $w\$$ of length at most $12^{(d-5)/4}+$ 1, recognizes whether $w\in L$.
An analogous result can be derived from Thm. 4.1 for fixed depth universal (i.e., shared parameter) transformers.
@@ 5 Log Depth Enables Graph Connectivity
In the graph connectivity problem (also referred to as STCON or the reachability problem), the input is a graph $G$, along with a source vertex $s$ and a target vertex $t$. The task is to determine if $G$ has a path from $s$ to $t$. This is a core problem at the heart of many computational questions in areas as diverse as network security, routing and navigation, chip design, and—perhaps most commonly for language models—multi-step reasoning. This problem is known to be complete for the class of logspace Turing machines (Reingold, 2008; Immerman, 1998), which means that, under common complexity conjectures, it cannot be solved accurately by fixed-depth transformers, which can only solve problems in the smaller class $\mathsf{TC}^{0}$. However, graph connectivity can be expressed by log-depth threshold circuits ($\mathsf{TC}^{1}$, Barrington & Maciel, 2000), which opens up a natural question: Can log-depth transformers, which are in $\mathsf{TC}^{1}$, solve graph connectivity? In Appendix C, we prove the following, which shows that the answer is yes:
Theorem 5.1 (Graph Connectivity).
@@
There exists a $(17,2,1)$-universal transformer $T$ with both causal and unmasked heads that, when unrolled $\lceil\log_{2}n\rceil$ times, solves connectivity on (directed or undirected) graphs over $n$ vertices: given the $n\times n$ adjacency matrix of a graph $G$, $n^{3}$ padding tokens, and $s,t\in\{1,\ldots n\}$ in unary notation, $T$ determines whether $G$ has a path from vertex $s$ to vertex $t$.
@@
Thus, while $\mathsf{NC}^{1}$ circuits (which have log depth) cannot express graph connectivity unless $\mathsf{NC}^{1}=\mathsf{NL}$, log-depth transformers can.
@@
Similar to the case of regular languages, this result also provides a concrete input length bound up to which a fixed-depth transformer can solve this problem.
@@
For any $d\in\mathbb{N}$, there exists a $d$-layer transformer with both causal and unmasked heads that solves connectivity on graphs with at most $2^{(d-18)/2}$ vertices.
@@ 6 Comparing Scaling Depth to Scaling Width or Chain of Thought
We now consider how increasing the depth of a transformer compares to other ways of adding more computation to the model. One natural option is to increase the width (i.e., model dimension; see Definition 2.1). Whereas slightly increasing depth expands expressive power beyond $\mathsf{TC}^{0}$, we show that achieving expressive power outside $\mathsf{TC}^{0}$ by only scaling width would require the width to grow superpolynomial in sequence length, which is infeasible. Another natural option is to expand the inference-time compute available to a pretrained model by adding chain-of-thought (CoT) steps. We now draw on related results in the literature to explore the efficiency of these alternatives.
@@
Our Corollaries 4.3 and 5.2 show that minimally growing a transformer’s depth allows it to express key problems that are likely outside $\mathsf{TC}^{0}$. In contrast, Thm. 6.1 shows that, if depth remains fixed, width must increase drastically with sequence length to enable expressive power outside $\mathsf{TC}^{0}$.
Theorem 6.1 (Width Scaling).
Let $T$ be a fixed-depth transformer whose width (model dimension) grows as a polynomial in $n$ and whose weights on input length $n$ (to accommodate growing width) are computable in $\mathsf{L}$. Then $T$ can be simulated in $\mathsf{L}$-uniform $\mathsf{TC}^{0}$.
Proof.
This extends Theorem 1 in Merrill & Sabharwal (2023a). For completeness, Appendix D gives a sketch. ∎
This result shows that, to solve reasoning problems outside $\mathsf{TC}^{0}$ over a context length $n$, growing depth is much more efficient than growing width. Of course, there may be other types of problems (e.g., those that are knowledge intensive or very parallelizable) where growing width might be more important than growing depth. Petty et al. (2024) provide an interesting empirical investigation of this choice on language modeling, semantic parsing, and other tasks.
@@
Merrill & Sabharwal (2024) analyze the power of transformers with $\mathrm{O}(\log n)$ CoT steps, showing it is at most $\mathsf{L}$. However, we have shown that transformers with $\Theta(\log n)$ depth can solve directed graph connectivity, which is $\mathsf{NL}$-complete: this suggests growing depth has some power beyond growing CoT unless $\mathsf{L}=\mathsf{NL}$. In fact, this can be extended (Li et al., 2024) to show transformers with $\mathrm{O}(\log n)$ CoT cannot solve any problem outside $\mathsf{TC}^{0}$.
Theorem 6.2 (CoT Scaling).
@@
Transformer with $\mathrm{O}(\log n)$ chain-of-thought steps recognize at most $\mathsf{L}$-uniform $\mathsf{TC}^{0}$.
Proof.
Follows from Li et al. (2024, Figure 10), which refines a previous upper bound of $\mathsf{L}$ (Merrill & Sabharwal, 2024). For completeness, Appendix D gives a sketch. ∎
@@
Thus, while giving a model $\mathrm{O}(\log n)$ CoT steps does not increase its expressive power beyond $\mathsf{TC}^{0}$, our Thms. 4.1 and 5.1 allow $\Theta(\log n)$ to solve key problems that are (likely) outside $\mathsf{TC}^{0}$. This demonstrates an advantage of dynamic depth over CoT as a form of inference-time compute for reasoning problems including regular language recognition and graph connectivity. It would be interesting to explore this comparison more generally for other problems.
7 Empirical Validation of Predicted Depth and Width Scaling
Our theory makes empirically testable predictions about the relationship between a model’s depth (and width) and the effective context length for key reasoning problems outside $\mathsf{TC}^{0}$. Specifically, as predicted by Thm. 4.1, is it the case that recognizing regular languages over strings of length $n$ empirically requires depth proportional to $\log n$? On the other hand, as predicted by Thm. 6.1, must the width scale as $\exp(\Theta(n))$ in practice in order to recognize strings of length $n$? Finally, if these relationships hold in practice, can we empirically quantify the constant factors?

### v2

@@ 4 Log Depth Enables Recognizing Regular Languages
Constant-depth transformers cannot recognize regular languages, a natural task closely related to state tracking (Liu et al., 2023; Merrill et al., 2024). Liu et al. (2023, Theorem 1) show that a non-uniform variant of log-depth transformer can recognize regular languages using an associative prefix-sum construction (cf. Hillis and Steele Jr, 1986; Blelloch, 1990). However, it is unclear whether standard transformers can implement their construction (Liu et al., 2023, Page 44) for two key reasons:
  1. 1.
    Their construction is non-uniform, in the sense that the parameters depend on the input length and depth of the network. This means that it is unclear whether a single transformer could implement it in a way that generalizes to inputs of any length.
  2. 2.
    Their result requires two strong simplifications to the transformer architecture: removing residual connections and assuming specific, nonstandard positional encodings. As discussed in Section 2.1, dealing with residual connections is particularly tricky in universal transformers, requiring proper memory management of cells in the residual stream so that outputs from the previous iteration of a layer do not interfere with a later iteration.
Our result, leveraging Lemma 1 (full proof in Appendix C), addressed both of these weaknesses. A single transformer with fixed parameters can recognize strings of any length; moreover this transformer can have residual connections and does not rely on specific positional encodings:
Theorem 1 (Regular Language Recognition).
@@
Theorem 1 thus reveals that running a transformer to $\Theta(\log n)$ depth on inputs of length $n$ unlocks new power compared to a fixed-depth transformer. If we do not care that the construction is uniform across layers, we can simplify the block of 8 layers that determines activeness to 1 layer: we simply hardcode the layer index $\ell$ and use a single transformer layer to compute $i\mod\ell$. Thus, the non-uniform construction results in a family of shallower transformers:
@@
These results can be extended beyond regular languages: if a $b$-layer transformer can perform some binary associative operation $\oplus:X\times X\to X$, then one can construct an $\Theta(b\log n)$ layer transformer that computes the iterated version on $n$ values, $x_{1}\oplus x_{2}\oplus\ldots\oplus x_{n}\in X$. One example is iterated matrix multiplication. For matrices from a fixed set (e.g., $k\times k$ boolean matrices), Theorem 1 already shows that this task can be performed. However, if the matrices are not from a fixed set (e.g., matrices over $\mathbb{Z}$ or $\mathbb{Q}$ or whose shape depends on $n$), then it is unclear whether log-depth transformers can solve the binary multiplication problem, and thus whether they can solve the iterated version.
Fixed Depth and Bounded Length Inputs.
Interestingly, while Theorems 1 and 1.1 are about log-depth transformers, they can be turned around to infer bounds on the input length up to which fixed depth transformers (i.e., depth fixed w.r.t. input length) can recognize regular languages. Specifically, given any regular language $L$ and a fixed $d$, Corollary 1.1 implies that there exists a depth $d$ transformer that can recognize strings $w\in L$ as long as $4\lceil\log_{2}\lvert w\rvert\rceil+5\leq d$, [^5] The inequality holds since Corollary 1.1 generalizes to show $T_{n}$ recognizes all strings of length $m\leq n$. which is satisfied if $4\;(1+\log_{2}\lvert w\rvert)+5\leq d$, i.e., $\lvert w\rvert\leq 2^{(d-9)/4}$:
@@
Let $L$ be a regular language over $\Sigma$ and \$$\$\not\in\Sigma$. For any $d\in\mathbb{N}$, there exists a causally masked $d$-layer transformer that, on any string $w\$$ of length at most $12^{(d-9)/4}+$ 1, recognizes whether $w\in L$.
An analogous result holds for universal (i.e., shared parameter) transformers from Theorem 1.
@@ 5 Log Depth Enables Graph Connectivity
In the graph connectivity problem (also referred to as STCON or the reachability problem), the input is a graph $G$, along with a source vertex $s$ and a target vertex $t$. The task is to determine if $G$ has a path from $s$ to $t$. This is a core problem at the heart of many computational questions in areas as diverse as network security, routing and navigation, chip design, and—perhaps most commonly for language models—multi-step reasoning. This problem is known to be complete for the class of logspace Turing machines (Reingold, 2008; Immerman, 1998), which means that, under common complexity conjectures, it cannot be solved accurately by fixed-depth transformers, which can only solve problems in the smaller class $\mathsf{TC}^{0}$. However, graph connectivity can be expressed by log-depth threshold circuits ($\mathsf{TC}^{1}$, Barrington and Maciel, 2000), which opens up a natural question: Can log-depth transformers, which are in $\mathsf{TC}^{1}$, solve graph connectivity? We prove the following, which shows that the answer is yes (see proof sketch below, full proof in Appendix D):
Theorem 2 (Graph Connectivity).
@@
There exists a $(17,2,1)$-universal transformer $T$ with both causal and unmasked heads that, when unrolled $\lceil\log_{2}n\rceil$ times, solves connectivity on (directed or undirected) graphs over $n$ vertices: given the $n\times n$ adjacency matrix of a graph $G$, $n^{3}$ padding tokens, and $s,t\in\{1,\ldots n\}$ in unary, $T$ checks whether $G$ has a path from vertex $s$ to vertex $t$.
Proof Sketch.
We will prove this for the more general case of a directed graph $G$ over $n$ vertices. Let $A\in\{0,1\}^{n\times n}$ be $G$’s adjacency matrix. The idea is to use the first $n^{2}$ tokens of the transformer to construct binary predicates $B_{\ell}(i,j)$ for $\ell\in\{0,1,\ldots,\lceil\log n\rceil\}$ capturing whether $G$ has a path of length at most $2^{\ell}$ from $i$ to $j$. To this end, the transformer will use the $n^{3}$ padding tokens to also construct intermediate ternary predicates $C_{\ell}(i,k,j)$ for $\ell\in\{1,\ldots,\lceil\log n\rceil\}$ capturing whether $G$ has paths of length at most $2^{\ell-1}$ from $i$ to $k$ and from $k$ to $j$. These two series of predicates are computed from each other iteratively, as in standard algorithms for graph connectivity:
$B_{0}(i,j)$ $\iff A(i,j)\ \vee\ i=j$ (1)
$C_{\ell+1}(i,k,j)$ $\iff B_{\ell}(i,k)\wedge B_{\ell}(k,j)$ (2)
$B_{\ell+1}(i,j)$ $\iff\exists k\;\textrm{s.t.}\;C_{\ell+1}(i,k,j)$ (3)
The crucial part is to construct a transformer that correctly operationalizes the computation of predicates $B_{\ell}$ and $C_{\ell}$. The input to the transformer is the adjacency matrix $A$ represented using $n^{2}$ tokens from $\{0,1\}$, followed by $n^{3}$ padding tokens $\square$, and finally the source and target nodes $s,t\in\{1,\ldots,n\}$ represented in unary notation using special tokens $a$ and $b$:
$A_{1,1}\ldots A_{1,n}\ A_{2,1}\ldots A_{2,n}\ \ldots\ldots\ A_{n,1}\ldots A_{n,n}\ \underbrace{\square\ldots\ldots\ldots\square}_{n^{3}}\ \underbrace{a\ldots\ldots a}_{s}\ \underbrace{b\ldots\ldots b}_{t}$
Initial Layers. The transformer starts off by using layer 1 to store $1/N,n,n^{2},s,$ and $t$ in the residual stream at every position. It then uses the next 15 layers to compute and store in the residual stream the semantic “coordinates” of each of the first $n^{2}+n^{3}$ token position, namely, $(i,j)$ for each of the first $n^{2}$ positions $p=in+j$ and $(i,k,j)$ for each of the next $n^{3}$ positions $p=n^{2}+(in^{2}+kn+j)$. Finally, Layer 17 of the transformer computes the predicate $B_{0}(i,j)$ at the first $n^{2}$ token positions.
Repeated Layers. The repeated layers alternates between computing the $C_{\ell}$ and the $B_{\ell}$ predicates for $\ell\in\{1,\ldots,\lceil\log n\rceil\}$. The idea is to compute $C_{\ell}(i,k,j)$ in the $n^{3}$ padding tokens by attending to positions $(i,k)$ and $(k,j)$ and retrieving $B_{\ell-1}(i,k)$ and $B_{\ell-1}(k,j)$. Similarly, $B_{\ell}(i,j)$ is computed in the $n^{2}$ input positions via uniform attention over padding positions $(i,k^{\prime},j)$ that store $C_{\ell}(i,k^{\prime},j)$.
Final Layers. Finally, in layer $182\lceil\log n\rceil+$ 18, the final token uses a head that attends with query $\langle\phi(s),\phi(t)\rangle$ corresponding to the source and target nodes $s$ and $t$ mentioned in the input, attending solely to the position with coordinates $(s,t)$, and retrieving the final value $B_{\lceil\log n\rceil}(s,t)$. ∎
@@
Thus, while $\mathsf{NC}^{1}$ circuits (which have log depth) cannot solve graph connectivity unless $\mathsf{NC}^{1}=\mathsf{NL}$, log-depth transformers can.
Fixed Depth and Bounded Length Inputs.
@@
As for regular languages, this result also provides a concrete input length bound up to which a fixed-depth transformer can solve this problem, namely when $18+2\lceil\log_{2}n\rceil\leq d$, which is satisfied if $18+2\;(1+\log_{2}n)\leq d$, i.e., $n\leq 2^{(d-20)/2}$:
@@
For any $d\in\mathbb{N}$, there exists a $d$-layer transformer with both causal and unmasked heads that, on any graph with at most $2^{(d-20)/2}$ vertices, solves the connectivity problem.
@@ 6 Comparing Scaling Depth to Scaling Width or Chain of Thought
Our results show that looping layers enables transformers to solve problems likely outside $\mathsf{TC}^{0}$. We now consider how looping compares in expressive power to other ways to add computation to transformers. Rather than increasing depth by repeating layers, one can increase a transformer’s width via a larger model dimension (Definition 1) or padding tokens (Pfau et al., 2024). Whereas slightly increasing depth likely expands expressive power beyond $\mathsf{TC}^{0}$, we show that achieving expressivity beyond $\mathsf{TC}^{0}$ via depth likely requires superpolynomial in width, which is intractable. In contrast to repeating layers, another way to extend inference-time computation is using chain-of-thought (CoT) steps. We thus compare the expressive power achieved repeated layers with CoT steps.
@@
Our Corollaries 1.2 and 2.1 show that minimally growing a transformer’s depth allows it to express key problems that are likely outside $\mathsf{TC}^{0}$. In contrast, Theorem 3 (which extends Merrill and Sabharwal (2023a); for completeness, Appendix E gives a sketch) shows that, if depth remains fixed, width must increase drastically with sequence length to enable expressive power outside $\mathsf{TC}^{0}$.
Theorem 3 (Width Scaling).
Let $T$ be a fixed-depth transformer whose width (model dimension or padding tokens; Pfau et al., 2024) grows as a polynomial in $n$ and whose weights on input length $n$ (to accommodate growing width) are computable in $\mathsf{L}$. Then $T$ can be simulated in $\mathsf{L}$-uniform $\mathsf{TC}^{0}$.
Thus, to solve reasoning problems outside $\mathsf{TC}^{0}$ over a context length $n$, growing depth is much more efficient than growing width. Of course, there may be other types of problems (e.g., those that are knowledge intensive or very parallelizable) where growing width might be more important than growing depth. Petty et al. (2024) provide an interesting empirical investigation of this choice on language modeling, semantic parsing, and other tasks.
@@
Merrill and Sabharwal (2024, Theorem 4) analyze the power of transformers with $\mathrm{O}(\log n)$ CoT steps, showing it is at most $\mathsf{L}$. However, we have shown that transformers with $\Theta(\log n)$ depth can solve directed graph connectivity, which is $\mathsf{NL}$-complete: this suggests growing depth has some power beyond growing CoT unless $\mathsf{L}=\mathsf{NL}$. In fact, the $\mathrm{O}(\log n)$ CoT steps result can be strengthened (Li et al., 2024, Figure 10; for completeness, Appendix E gives a sketch) to an upper bound of $\mathsf{TC}^{0}$:
Theorem 4 (CoT Scaling).
@@
Transformers with $\mathrm{O}(\log n)$ chain-of-thought steps can only recognize languages in $\mathsf{L}$-uniform $\mathsf{TC}^{0}$.
@@
Thus, while giving a model $\mathrm{O}(\log n)$ CoT steps does not increase its expressive power beyond $\mathsf{TC}^{0}$, our Theorems 1 and 2 allow $\Theta(\log n)$ to solve key problems that are (likely) outside $\mathsf{TC}^{0}$. This demonstrates an advantage of dynamic depth over CoT as a form of inference-time compute for reasoning problems including regular language recognition and graph connectivity. It would be interesting to explore this comparison more generally for other problems.
7 Experiments: Learning to Recognize Regular Languages
Our theory characterizes the depth and width required to express regular language recognition and graph connectivity. Specifically, Theorem 1 predicts that recognizing regular languages over strings of length $n$ is empirically possible with depth proportional to $\log n$. On the other hand, Theorem 3 predicts that the width would need to scale superpolynomially. Here, we aim to empirically measure how much depth and width transformers require in practice when trained to recognize regular languages. We will find that expressibility and learnability are highly aligned here: transformers with log depth can learn to recognize regular languages, whereas depth must increase exponentially with $n$. Moreover, we can empirically quantify the constant factors in these relationships.